The XYZ Company produces a single product: very large, hand stitched national flags for the Peoples’ Liberation Army.   Each year XYZ sells around 40,000 flags to the PLA.

In 20×2, XYZ had budgeted production of 40,000 flags. Each flag requires 2.8 square meters of cloth, the only direct material. The budgeted cost of the cloth is $1.20 per square meter. All emblems on the flag are stitched by hand, a time consuming process. Budgeted direct labor time for each flag is 4.2 hours. The budgeted cost of direct labor is $16/hour.

Manufacturing Overhead (MNFG OVHD) is allocated to the product based on machine hours. 0.8 machine hours are budgeted for the production of each flag. Variable MNFG OVHD is budgeted at $3.60 per machine hour. Fixed MNFG OVHD is budgeted at $35,200 for the year.

During 20×2, XYZ produced 38,500 flags. They purchased 116,000 square meters of cloth at a total cost of $141,520 and they used 109,500 square meters in production. XYZ used 160,500 direct labor hours at a total cost of $2,425,000 and 36,000 machine hours. The actual Variable MNFG OVHD costs incurred were $114,000 and the actual Fixed MNFG OVHD costs incurred were $36,120.

Required: Compute each of the following: (3 marks each)

Note, the direct material price variance is recognized at the time of purchase. Also, for each variance, you must indicate the AMOUNT and DIRECTION (F or U) of the variance.

1. The Direct Material Price Variance

2. The direct Material Efficiency Variance

3. The Direct Labor Price Variance

4. The Direct Labor Efficiency Variance

5. The Variable MNFG OVHD Spending Variance

6. The Variable MNFG OVHD Efficiency Variance  

7. The Fixed MNFG OVHD Spending Variance  

8. The Fixed MNFG OVHD production Volume Variance   

9. Present the general journal entry (in good form) needed to record the actual Fixed MNFG OVHD Costs incurred during 20×2.   

10. Present the General Journal Entry needed to record the allocation of Fixed MNFG OVHD to the product during 20×2.

11. Present the General Journal Entry needed at the end of 20×2 to close the Fixed MNFG OVHD accounts                
Answer #1
1 Direct material price variance=AQ*(SR-AR)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
AQ=Actual material purchased=116000 Sq. mtr.
SR=Standard Rate=$1.20 per sq mtr
AR=Actual rate=Total cost of material purchased/AQ=141520/116000=$ 1.22 per sq mtr
Direct material price variance=116000*(1.20-1.22)=-2320=$ 2320 Unavorable
2 Direct material efficiency variance=SR*(SQ-AQ)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
SR=Standard Rate=$1.20 per sq mtr
SQ=Standard Quantity=Actual units produced*Material required per unit=38500*2.8=107800 sq mtr
AQ=Actual quantity used=109500 sq mtr
Direct material efficiency variance=1.20*(107800-109500)=-2040=$ 2040 Unfavorable
3 Direct labor price variance=ALH*(SR-AR)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
ALH=Actual labor hours=160500 Hours
SR=Standard labor rate per hour=$16
AR=Actual labor rate per hour=Actual labor cost/ALH=2425000/160500=$ 15.11
Direct labor price variance=160500*(16-15.11)=$ 142845 Favorable
4 Direct labor efficiency variance=SR*(SLH-ALH)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
SR=Standard labor rate per hour=$16
SLH=Standard labor hours=Actual units produced*Labor hours per unit=38500*4.2=161700 hours
ALH=Actual labor hours=160500 Hours
Direct labor efficiency variance=16*(161700-160500)=19200=$ 19200 Favorable
5 Variable MNFG OVHD spending variance=AMH*(SR-AR)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
AMH=Actual machine hours=36000 Hours
SR=Standard variable overhead rate=$3.60 per machine hour
AR=Actual variable overhead rate=Actual variable overhead/AMH=114000/36000=$ 3.17 per machine hour
Variable Variable MNFG OVHD spending variance=36000*(3.60-3.17)=15480=$ 15480 Favorable
6 Variable MNFG OVHD efficiency variance=SR*(SMH-AMH)
If the answer is positive,variance is favorable.Otherwise,unfavorable
Where
SR=Standard variable overhead rate=$3.60 per machine hour
SMH=Standard machine hours=Actual units produced*Standard machine hours per unit=38500*0.8=30800 machine hours
AMH=Actual machine hours=36000 Hours
Variable MNFG OVHD efficiency variance=3.60*(30800-36000)=-18720=$ 18720 Unfavorable